OpenStudy (anonymous):

in a triangle a the longer leg is 4 more than the shorter leg and the hypotenuse is 4 more than the longer leg. Find the Short Leg=_______ Long Leg=_______ Hypotenuse=_______

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Let s be the length of the shorter leg. Then from the problem narrative,\[(s+4+4)^2=s^2+(s+4)^2 \]12^2Solve the above for s. s = 12 {s, s + 4, s + 8} {12, 16, 20} Check out the above for any mistakes.

OpenStudy (anonymous):

is that a specific formula for this type of problem?

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

The problem text indicates a right triangle. For a right triangle, the square of the hypotenuse is equal to the sum of the squares upon the legs. The answer is 12, 16 and 20 in that order.